Scores on the quantitative portion of an exam have a mean of

586

and a standard deviation of

140.

Assume the scores are normally distributed. What percentage of students taking the quantitative exam score above

621?

What percentage of students taking the quantitative exam score above

621​?

​(Round to the nearest whole number as​ needed.)
Answer #1

Solution :

Given that,

mean = = 586

standard deviation = = 140

P(x >621 ) = 1 - P(x<621 )

= 1 - P[(x -) / < (621-586) / 140]

= 1 - P(z <0.25 )

Using z table

= 1 - 0.5987

= 0.4013

=40.13%

=40% rounded